• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name A. Camsonne

    Log entry time 15:56:08 on April25,2003

    Entry number 98572

    keyword=Day shift summary

    No beam expected today because of the open house

    8:40 turned target lasers off for survey and work on the target

    9:00 Septum IOC rebooted

    9:30 Put reference cell to manual control so that people can control it from the hall

    10:20 Vince turn the lasers on

    11:37 Septum IOC rebooted twice

    12:50 Current on septum magnet
    Paul Brindza put some current in the septum magnet

    14:53 Septum IOC rebooted by itself

    15:20 Network lost with IOCs
    Called MCC who pointed me to software on Call
    problem with gateway : 129.57.255.4
    After delogging and relogging hascsbc2 everything looks fine
    I hope this won't happen again while we are running
    Maybe some work going on the network and IOCs

    16:00 Karl takes over